# Theory and Interpretability of Quantum Extreme Learning Machines: a Pauli-transfer matrix approach

### v1

@@ Abstract
Quantum reservoir computers (QRCs) have emerged as a promising approach to quantum machine learning, since they utilize the natural dynamics of quantum systems for data processing and are simple to train. Here, we consider $n$-qubit quantum extreme learning machines (QELMs) with continuous-time reservoir dynamics. QELMs are memoryless QRCs capable of various ML tasks, including image classification and time series forecasting. We apply the Pauli transfer matrix (PTM) formalism to theoretically analyze the influence of encoding, reservoir dynamics, and measurement operations, including temporal multiplexing, on the QELM performance. This formalism makes explicit that the encoding determines the complete set of (nonlinear) features available to the QELM, while the quantum channels linearly transform these features before they are probed by the chosen measurement operators. Optimizing a QELM can therefore be cast as a decoding problem in which one shapes the channel-induced transformations such that task-relevant features become available to the regressor. The PTM formalism allows one to identify the classical representation of a QELM and thereby guide its design towards a given training objective. As a specific application, we focus on learning nonlinear dynamical systems and show that a QELM trained on such trajectories learns a surrogate-approximation to the underlying flow map.
@@ I Introduction
QRCs have enjoyed significant attention since their introduction [1], featuring applications in various domains, such as image classification [2, 3, 4, 5], time series forecasting [1, 6, 7, 8, 9, 10], fluid dynamics [11, 12], or post-processing of quantum experiments [13, 14, 15, 16]. A major advantage of QRCs is their suitability for noisy intermediate-scale quantum (NISQ) hardware, primarily because the training process is restricted to the classical readout layer, thereby rendering the training problem convex and avoiding the barren plateau problem associated with variational quantum algorithms [17]. QRCs are often resilient to, and can even benefit from, environmental noise [18, 19, 20, 21]. The relevance of nonlinear input transformations has been analyzed for QRCs in [22, 23, 24], and universality conditions have been derived in [25, 26, 27]. Notably, QRCs are closely related to quantum kernel methods [28]—a fact that can be utilized to find optimal measurement operators [29] and harness insights from the classical domain [30, 31, 32]. For general overviews on QRCs, we refer to [33, 34].
@@
A specific QRC variant without any coherently evolving internal state is the QELM [35, 33, 34, 2]. Its classical counterpart, the ‘extreme learning machine’, is essentially a feedforward network with fixed internal architecture [36], similarly to a random-feature model [37]. A QELM differs from a conventional quantum neural network with fixed internal parameters in two important ways: first, instead of gate-based architectures, it typically employs the natural dynamics of a physical quantum system (a reservoir). Second, instead of optimizing a measurement output for a certain task, one collects multiple unspecific measurements into a large readout vector and feeds it into a classical regression or classification algorithm. The main task in optimizing QELMs thus consists in tuning the quantum properties of the reservoir to maximize information processing and being able to properly decode relevant features from the measurements [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 14]. Identifying possible quantum advantages, e.g., related to the exponentially large Hilbert space into which data is encoded, is an active research topic [48, 34, 49, 41, 3, 45]. However, the exponential concentration effect emerging for large qubit numbers is a phenomenon closely related to barren plateaus [17, 50] that can strongly reduce QELM performance [51].
@@
The present work aims at a interpretable framework that can explain why and which failures in QELMs occur and thereby support systematic design of QELMs. To this end, we will also analyze well-known optimization tools like temporal multiplexing [1] or information processing capacity [52, 53] in the light of the PTM formalism. We use chaotic dynamical systems as benchmarks since they have a known ground-truth model and are in principle discoverable from training data [54]. Our findings also hold implications for general QML models.
@@ II Model
We briefly review the formalism and specific implementation of a QELM used in this work [1, 33, 34]. Let $\mathcal{M}(\mathcal{H})$ denote the space of $d^{2}$-dimensional Hermitian matrices on the $d$-dimensional Hilbert space $\mathcal{H}$, where $d=2^{n}$ in terms of the number of qubits $n$. Given classical inputs $\mathbf{x}\in\mathbb{R}^{D}$, a QELM is defined by the output function
$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})]$ (1)
@@
$\min_{\mathbf{w}\in\mathbb{R}^{m}}\sum_{i=1}^{P}\mathcal{L}\left(f(\mathbf{x}_{i}),\,y_{i}\right)+\frac{\lambda}{2}\|\mathbf{w}\|^{2}_{2}$ (2)
@@
$\mathbf{w}^{*}=(\boldsymbol{\Psi}\boldsymbol{\Psi}^{\top}+\lambda\mathbbm{1}_{m})^{-1}\boldsymbol{\Psi}\mathbf{y},\qquad\Psi_{ki}=\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i}))$ (3)
@@
When optimizing over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d}$, the QELM takes the form of a quantum kernel model [28, 29] (see Appendix A for a summary). In the present work, we do not use that formalism, but instead work with a preselected observable set and optimize only over the associated weights $w_{k}$.
@@ II.1 Feature map
Here, $S(\mathbf{x})$ is a data encoding unitary, $U_{R}$ is a reservoir unitary, and the initial state $|\Phi_{0}\rangle$ is taken to be the ground state $|\Phi_{0}\rangle=|0\rangle^{\otimes n}$. The unitary $U_{R}(t)=\exp(-\mathrm{i}tH_{R})$ is defined by time evolution under a reservoir Hamiltonian $H_{R}$. Specifically, we consider Haar-random Hamiltonians $H_{R}$ as well as two variants of a transverse field Ising model (TFIM)
@@ II.2 Encoding schemes
We consider (parallel) product state encodings with $D=n$, producing the state
$\rho_{\text{enc}}(\mathbf{x})=S(\mathbf{x})\rho_{0}S(\mathbf{x})^{\dagger}=\bigotimes_{j=1}^{n}|\psi_{\text{enc}}(x_{j})\rangle\langle\psi_{\text{enc}}(x_{j})|,\qquad S(\mathbf{x})=\bigotimes_{j=1}^{n}S_{j}(x_{j})$ (7)
where the unitary $S_{j}$ act on the $j$-th qubit. Specifically, in our numerical experiments we mostly work with amplitude encoding [55]. This scheme comes in two variants, which lead to different expressivities: one encodes data on half a great circle of the Bloch sphere
@@
and one on the full great circle
@@
In a few cases, we also consider rotational encoding [55], defined by the unitary
$S_{j}(x_{j})=\exp(-i\pi x_{j}\sigma^{p_{j}})=R_{p_{j}}(2\pi x),\qquad x\in[0,1],$ (10)
@@ III Theory
QML models, including and QELMs, have often been analyzed from the Fourier perspective, which is natural for Hamiltonian-based encoding [48, 51, 49]. In the following, we instead employ the Pauli transfer matrix (PTM) formalism, which allows a more direct identification of the relevant features [33, 43, 56, 57, 58] and has been previously used for the analysis of QRCs [33, 25, 46].
@@ III.1 Pauli-transfer formulation of a QELM
giving rise to a feature vector $\boldsymbol{\upphi}(\mathbf{x})\in\mathbb{R}^{d^{2}}$. For product state encodings, the components of $\phi_{k}(x)$ correspond to all possible $d^{2}$ monomial combinations of the single-Pauli features (see Section III.2). Any quantum channel $\mathcal{E}$ acts linearly on this vector via the Pauli transfer matrix:
@@
Consider the QELM in Eq. 1 with a subset of measured Pauli operators $\{P_{k}\}_{k=1}^{{m}}$, ${m}\leq d^{2}$. This selection can be represented by a matrix $\mathcal{S}\in\mathbb{R}^{{m}\times d^{2}}$ (a subset of rows of the identity matrix) that extracts the corresponding rows from $T_{\mathcal{E}}$. The readout vector $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{{m}}$ contains their expectation values $F_{k}(\mathbf{x})=\mathrm{tr}\bigl[P_{k}\,\mathcal{E}(\rho_{\text{enc}}(\mathbf{x}))\bigr]$. Expanding $\rho_{\text{enc}}$ as in Eq. 13 and using Eq. 14 gives the model output function
$f(\mathbf{x})=\mathbf{w}^{\top}\,\mathbf{F}(\mathbf{x})=\mathbf{w}^{\top}\,T_{\mathcal{E}}^{({m})}\,\boldsymbol{\upphi}(\mathbf{x})=\sum_{k=1}^{m}w_{k}\sum_{j=0}^{d^{2}-1}(\mathcal{S}T_{\mathcal{E}})_{kj}\phi_{j}(\mathbf{x}),$ (16)
where $\mathbf{w}\in\mathbb{R}^{{m}}$ are classical weights and $T_{\mathcal{E}}^{({m})}\equiv\mathcal{S}T_{\mathcal{E}}$ is the ${m}\times d^{2}$ block of $T_{\mathcal{E}}$ restricted to the measured observables. This formulation explicitly shows that $f(\mathbf{x})$ is a linear function of the features $\boldsymbol{\upphi}(\mathbf{x})$ generated solely by the encoding and mixed by the reservoir channel $T_{\mathcal{E}}$.
@@
$(T_{\mathcal{E}}^{({m})})_{kj}=\frac{1}{d}\sum_{\ell=1}^{\mu}\mathrm{tr}\bigl(P_{k}\,K_{\ell}P_{j}K_{\ell}^{\dagger}\bigr)=\mathrm{tr}\bigl(P_{k}^{(H)}P_{j}\bigr)$ (17)
@@ III.2 Feature encoding
with $\phi_{0}^{(j)}=\mathrm{tr}(\rho_{j})=1$ (normalization), $\sigma^{0}\equiv\mathbbm{1}$ and $\sigma^{x,y,z}\equiv X,Y,Z$ as before. The three non-trivial components of $\boldsymbol{\upphi}^{(j)}$ define the Bloch vector $\boldsymbol{\varphi}^{(j)}=\{\phi_{x}^{(j)},\phi_{y}^{(j)},\phi_{z}^{(j)}\}$, which fulfills $\|\boldsymbol{\varphi}^{(j)}\|^{2}=1$ for pure states (mainly considered here) and $\|\boldsymbol{\varphi}^{(j)}\|^{2}\leq 1$ for mixed states. Since both the state and the Pauli basis factorize across qubits, the $n$-qubit feature vector is simply
@@
The subscript $a_{1}\ldots a_{n}$ corresponds to the explicit site representation of the index $k$ in Eq. 13. The $\phi_{k}$ thus run over all possible combinations of the single-qubit features in Eq. 21: for each qubit one chooses one of the four factors $\{1,\phi^{(j)}_{x},\phi^{(j)}_{y},\phi^{(j)}_{z}\}$, yielding a total of at most $d^{2}=4^{n}$ features. If the measured observables are Pauli strings of product form $P=\bigotimes_{j=1}^{n}\sigma^{a_{j}}$, the corresponding readout feature factorizes analogously as $\phi^{\prime}_{a_{1}\ldots a_{n}}(\mathbf{x})=\prod_{j=1}^{n}\bigl(\boldsymbol{\upphi}^{{}^{\prime}(j)}(x_{j})\bigr)_{a_{j}}$ with the post-reservoir feature vector $\boldsymbol{\upphi}^{\prime}$. The readout features resulting from some simple encoding schemes are summarized in Table 1. Effects of feature mixing by a unitary channel are discussed in the next section.
@@
Factor in observable $P$ Amplitude on $[0,1]$ Rotational
$\cdots\otimes X_{k}\otimes\cdots$ $2\sqrt{(1-u_{k})u_{k}}$ $\sin(\pi u_{k})$
$\cdots\otimes Y_{k}\otimes\cdots$ 0 0
$\cdots\otimes Z_{k}\otimes\cdots$ $2u_{k}-1$ $\cos(\pi u_{k})$
$\cdots\otimes\tfrac{1}{2}(\mathbb{I}+Z_{k})\otimes\cdots$ $u_{k}$ $\cos^{2}(\pi u_{k}/2)$
Table 1: Pauli–basis projections $\langle\psi_{\text{enc}}|P|\psi_{\text{enc}}\rangle$ of an encoded product state $\rho_{\text{enc}}(\mathbf{u})$ for rotational encoding (10) and amplitude encoding (8). Here $P=\bigotimes_{k=1}^{n}\sigma_{a_{k}}$ denotes a Pauli operator ($\sigma_{a_{k}}\in\{\mathbb{I},X,Y,Z\}$) and the classical datum entering qubit $k$ is $u_{k}$. Replacing $u_{k}\!\to\!u_{k}^{2}$ transforms the first column to the amplitude encoding variant (9). For convenience, we have included the projection operator $\frac{1}{2}(\mathbb{I}+Z_{k})$. One observes that each Pauli factor $\sigma_{a_{k}}$ (first column) in an observable $P$ produces the corresponding single-qubit feature $\phi_{a_{k}}^{(k)}(u_{k}):=\mathrm{tr}\!\big(\sigma_{a_{k}}\,\rho_{k}(u_{k})\big)$ (2nd and 3rd column). For instance, the $X_{k}$ row gives $\phi_{x}^{(k)}(u_{k})$ and the $Z_{k}$ row gives $\phi_{z}^{(k)}(u_{k})$.
IV PTM analysis for general quantum channels
For separable quantum channels, the PTM factorizes as $T=\bigotimes_{j=1}^{n}T_{j}$ into single-qubit PTMs $T_{j}$. The action of a local unitary is discussed in Section D.1. To understand and optimize a QELM with general nonlocal quantum channels, we must analyze the full PTM [Eqs. 16 and 17]. The optimization can be split into several subtasks that define the fundamental reconstruction problem in (quantum) reservoir computing [42, 46, 59, 14]:

### v2

@@ Abstract
Quantum reservoir computers (QRCs) have emerged as a promising approach to quantum machine learning, since they utilize the natural dynamics of quantum systems for data processing and are simple to train. Here, we consider $n$-qubit quantum extreme learning machines (QELMs) with initial-state encoding and continuous-time reservoir dynamics. QELMs are memoryless QRCs capable of various ML tasks, such as image classification and time series forecasting. We apply the Pauli transfer matrix (PTM) formalism to theoretically analyze the influence of encoding, reservoir dynamics, and measurement operations (including temporal multiplexing) on the QELM performance. This formalism makes explicit that the encoding determines the complete set of (nonlinear) features available to this type of QELM, while the quantum channels linearly transform these features before they are probed by the chosen measurement operators. Optimizing such a QELM can therefore be cast as a decoding problem in which one shapes the channel-induced transformations such that task-relevant features become available to the regressor, effectively reversing the information scrambling of a unitary. We show how operator spreading under unitary evolution determines feature decodability, which underlies the nonlinear processing capacity of the reservoir. The PTM formalism yields a nonlinear vector (auto-)regression model as the classical representation of a QELM. As a specific application, we focus on learning nonlinear dynamical systems and show that a QELM trained on such trajectories learns a surrogate-approximation to the underlying flow map.
@@ I Introduction
QRCs have enjoyed significant attention since their introduction [1], featuring applications in various domains, such as image classification [2, 3, 4, 5], time series forecasting [1, 6, 7, 8, 9, 10], fluid dynamics [11, 12], or post-processing of quantum experiments [13, 14, 15, 16]. A major advantage of QRCs is their suitability for noisy intermediate-scale quantum (NISQ) hardware, primarily because the training process is restricted to the classical readout layer, thereby rendering the training problem convex and avoiding the barren plateau problem associated with variational quantum algorithms [17, 18]. QRCs are often resilient to, and can even benefit from, environmental noise [19, 20, 21, 22]. The relevance of nonlinear input transformations has been analyzed for QRCs in [23, 24, 25], and universality conditions have been derived in [26, 6, 27, 28, 29, 30]. Notably, QRCs are closely related to quantum kernel methods [31]—a fact that can be utilized to find optimal measurement operators [32] and harness insights from the classical domain [33, 34, 35]. For general overviews on QRCs, we refer to [36, 37].
@@
A specific QRC variant without any coherently evolving internal state is the QELM [38, 36, 37, 2]. Its classical counterpart, the ‘extreme learning machine’, is essentially a feedforward network with fixed internal architecture [39], similarly to a random-feature model [40]. A QELM differs from a conventional quantum neural network with fixed internal parameters in two important ways: first, instead of gate-based architectures, it typically employs the natural dynamics of a physical quantum system (a reservoir). Second, instead of optimizing a measurement output for a certain task, one usually collects multiple unspecific measurements into a large readout vector and feeds it into a classical regression or classification algorithm. The main task in optimizing QELMs thus consists in tuning the quantum properties of the reservoir to maximize information processing and being able to properly decode relevant features from the measurements [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 14]. Identifying possible quantum advantages, e.g., related to the exponentially large Hilbert space into which data is encoded, is an active research topic [51, 37, 52, 44, 3, 48]. However, the exponential concentration effect emerging for large qubit numbers is a phenomenon conceptually related to barren plateaus and can strongly reduce QELM performance [53].
@@
The present study aims at an interpretable framework that can explain why and which failures in QELMs occur and thereby support their systematic design. To this end, we will also analyze well-known optimization tools like temporal multiplexing [1] or information processing capacity [54, 55] in the light of the PTM formalism. We use chaotic dynamical systems as benchmarks since they have a known ground-truth model and are in principle discoverable from training data [56]. The developed formalism is restricted to $n$-qubit QELMs in which the encoding is time-independent and separated from reservoir evolution, hence excluding models with dynamical input injection [57, 14] or generic continuous-variable (infinite-dimensional) encodings [58, 59, 24]. Our findings also hold implications for other QML models that fall into this category.
@@ II Model
We briefly present the formalism of QELMs and the specific implementation used in this work [1, 36, 37]. Let $\mathcal{M}(\mathcal{H})$ denote the space of $d\times d$ Hermitian matrices (real vector space of dimension $d^{2}$) on the $d$-dimensional Hilbert space $\mathcal{H}$, where $d=2^{n}$ in terms of the number of qubits $n$. Given classical inputs $\mathbf{x}\in\mathbb{R}^{D}$, a QELM is defined by the output function
$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})],$ (1)
@@
$\min_{\mathbf{w}\in\mathbb{R}^{m}}\sum_{i=1}^{P}\mathcal{L}\left(f(\mathbf{x}_{i}),\,y_{i}\right)+\frac{\lambda}{2}\|\mathbf{w}\|^{2}_{2},$ (2)
@@
$\hat{\mathbf{w}}=(\boldsymbol{\Psi}\boldsymbol{\Psi}^{\top}+\lambda\mathbbm{1}_{m})^{-1}\boldsymbol{\Psi}\mathbf{y},\qquad\Psi_{ki}=\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i})),$ (3)
@@
When optimizing over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d^{2}}$, the QELM takes the form of a quantum kernel model [31, 32] (see Appendix A for a summary). In the present work, we do not use that formalism, but instead work with a preselected observable set and optimize only over the associated weights $w_{k}$.
@@ II.1 Feature map
Here, $S(\mathbf{x})$ is a data encoding unitary, $U_{R}$ is a reservoir unitary, and we take the initial state as $|\Phi_{0}\rangle=|0\rangle^{\otimes n}$. The unitary $U_{R}(t)=\exp(-\mathrm{i}tH_{R})$ is defined by time evolution under a reservoir Hamiltonian $H_{R}$. Specifically, we consider random Hamiltonians $H_{R}$ as well as two variants of a transverse field Ising model (TFIM)
@@ II.2 Encoding schemes
We consider time-independent (parallel) product state encodings with $D=n$, producing the state
$\rho_{\text{enc}}(\mathbf{x})=S(\mathbf{x})\rho_{0}S(\mathbf{x})^{\dagger}=\bigotimes_{j=1}^{n}|\psi_{\text{enc}}(x_{j})\rangle\langle\psi_{\text{enc}}(x_{j})|,\qquad S(\mathbf{x})=\bigotimes_{j=1}^{n}S_{j}(x_{j}),$ (7)
where the unitary $S_{j}$ act on the $j$-th qubit. In our numerical experiments we employ amplitude and rotational encoding schemes [60]. The former comes in two variants, where one either encodes data on half a great circle of the Bloch sphere:
@@
or on the full great circle:
@@
Rotational encoding is defined by the unitary
$S_{j}(x_{j})=\exp(-i\pi x_{j}\sigma^{p_{j}})=R_{p_{j}}(2\pi x_{j}),\qquad x\in[0,1],$ (10)
@@ III Theory
QML models, including QELMs, have often been analyzed from the Fourier perspective, which is natural for Hamiltonian-based encoding [51, 53, 52]. In the following, we instead employ the Pauli transfer matrix (PTM) formalism, which allows a more direct identification of the relevant features for models with initial-state encoding [36, 46, 61, 62, 63] and has been previously used for the analysis of QRCs [36, 28, 49].
@@ III.1 Pauli-transfer formulation of a QELM
giving rise to a feature vector $\boldsymbol{\upphi}(\mathbf{x})\in\mathbb{R}^{d^{2}}$. For product state encodings, the components of $\phi_{k}(\mathbf{x})$ correspond to all possible $d^{2}$ monomial combinations of the single-Pauli features (see Section III.2). Any quantum channel $\mathcal{E}$ acts linearly on this vector via the Pauli transfer matrix:
@@
Consider the QELM in Eq. 1 with a subset $\mathcal{S}$ of measured Pauli operators $\{P_{k}\}_{k\in\mathcal{S}}$, ${m}\equiv|\mathcal{S}|\leq d^{2}$. This selection can be represented by a matrix $S\in\mathbb{R}^{{m}\times d^{2}}$ (a subset of rows of the identity matrix) that extracts the corresponding rows from $T_{\mathcal{E}}$. The readout vector $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{{m}}$ contains their expectation values $F_{k}(\mathbf{x})=\mathrm{tr}\bigl[P_{k}\,\mathcal{E}(\rho_{\text{enc}}(\mathbf{x}))\bigr]$. Expanding $\rho_{\text{enc}}$ as in Eq. 13 and using Eq. 14 gives the model output function
$f(\mathbf{x})=\mathbf{w}^{\top}\,\mathbf{F}(\mathbf{x})=\mathbf{w}^{\top}\,T_{\mathcal{E}}^{({m})}\,\boldsymbol{\upphi}(\mathbf{x})=\sum_{k\in\mathcal{S}}w_{k}\sum_{j=0}^{d^{2}-1}(ST_{\mathcal{E}})_{kj}\phi_{j}(\mathbf{x}),$ (16)
where $\mathbf{w}\in\mathbb{R}^{{m}}$ are classical weights and $T_{\mathcal{E}}^{({m})}\equiv ST_{\mathcal{E}}$ is the ${m}\times d^{2}$ block of $T_{\mathcal{E}}$ restricted to the measured observables. This formulation explicitly shows that $f(\mathbf{x})$ is a linear function of the features $\boldsymbol{\upphi}(\mathbf{x})$ generated solely by the encoding and mixed by the reservoir channel $T_{\mathcal{E}}$.
@@
$(T_{\mathcal{E}}^{({m})})_{kj}=\frac{1}{d}\sum_{\ell=1}^{\mu}\mathrm{tr}\bigl(P_{k}\,K_{\ell}P_{j}K_{\ell}^{\dagger}\bigr)=\frac{1}{d}\mathrm{tr}\bigl(P_{k}^{(H)}P_{j}\bigr),$ (17)
@@ III.2 Feature encoding
with $\phi_{0}^{(j)}=\mathrm{tr}(\rho_{j})=1$ (normalization), $\sigma^{0}\equiv\mathbbm{1}$ and $\sigma^{x,y,z}\equiv X,Y,Z$ as before. The three non-trivial components of $\boldsymbol{\upphi}^{(j)}$ define the Bloch vector, which has unit magnitude for pure states (see Appendix D for further discussion). Since both the state and the Pauli basis factorize across qubits, the $n$-qubit feature vector is
@@
The subscript $a_{1}\ldots a_{n}$ corresponds to the explicit site representation of the index $k$ in Eq. 13. The $\phi_{k}$ thus run over all possible combinations of the single-qubit features in Eq. 21: for each qubit one chooses one of the four factors $\{1,\phi^{(j)}_{x},\phi^{(j)}_{y},\phi^{(j)}_{z}\}$, yielding a total of at most $d^{2}=4^{n}$ features. The readout features resulting from some simple encoding schemes are summarized in Table 1. Effects of feature mixing by a unitary channel are discussed in the next section.
@@
Factor in observable $P$ Amplitude on $[0,1]$ Amplitude on [1,1][-1,1] Rotational
$\cdots\otimes X_{k}\otimes\cdots$ $2\sqrt{(1-u_{k})u_{k}}$ $2u_{k}\sqrt{1-u_{k}^{2}}$ $\sin(2\pi u_{k})$
$\cdots\otimes Y_{k}\otimes\cdots$ 0 0 0
$\cdots\otimes Z_{k}\otimes\cdots$ $2u_{k}-1$ $2u_{k}^{2}-1$ $\cos(2\pi u_{k})$
$\cdots\otimes\tfrac{1}{2}(\mathbb{I}+Z_{k})\otimes\cdots$ $u_{k}$ $u_{k}^{2}$ $\cos^{2}(\pi u_{k})$
Table 1: Pauli–basis projections $\langle\psi_{\text{enc}}|P|\psi_{\text{enc}}\rangle$ of an encoded product state $\rho_{\text{enc}}(\mathbf{u})$ for amplitude encoding (8) or (9) as well as rotational encoding (10). Here $P=\bigotimes_{k=1}^{n}\sigma_{a_{k}}$ denotes a Pauli operator ($\sigma_{a_{k}}\in\{\mathbb{I},X,Y,Z\}$) and the classical datum entering qubit $k$ is $u_{k}$. For convenience, we have included the projection operator $\frac{1}{2}(\mathbb{I}+Z_{k})$. One observes that each Pauli factor $\sigma_{a_{k}}$ (1st column) in an observable $P$ produces the corresponding single-qubit feature $\phi_{a_{k}}^{(k)}(u_{k}):=\mathrm{tr}\!\big(\sigma_{a_{k}}\,\rho_{k}(u_{k})\big)$ (2nd–4th columns). For instance, the $X_{k}$ row gives $\phi_{x}^{(k)}(u_{k})$ and the $Z_{k}$ row gives $\phi_{z}^{(k)}(u_{k})$.
IV PTM analysis for a unitary quantum channel
For separable (product) quantum channels, the PTM factorizes as $T=\bigotimes_{j=1}^{n}T_{j}$ into single-qubit PTMs $T_{j}$. The action of a local unitary is discussed in Section D.1. To understand and optimize a QELM with general nonlocal quantum channels, we must analyze the full PTM [Eqs. 16 and 17]. We focus in the following on the unitary case. The optimization can be split into several subtasks that define the fundamental reconstruction problem in (quantum) reservoir computing with explicit feature maps [45, 49, 64, 14]: